OpenStudy (anonymous):

solve the equation and show the steps 16+x^2-8x=0

OpenStudy (campbell_st):

it can be written x^2 - 8x + 16 = 0 find the factors of 16 that add to -8.... hint both are negative

OpenStudy (anonymous):

the factors would be -4 and -4 but what do i do with that?

OpenStudy (campbell_st):

then you have x-4)(x -4) = 0 solve x - 4 = 0 this will be the solution

OpenStudy (anonymous):

would i use foil?

OpenStudy (anonymous):

would the answer be x=4?

OpenStudy (campbell_st):

no... you have (x -4)(x-4)=0 either expression can be zero to give an zero answer 0*(x-4) = 0 or (x -4)*0 = 0 so what value of x makes x - 4= 0

OpenStudy (campbell_st):

thats the answer x = 4 its called repeated root.

OpenStudy (anonymous):

oooo ok i get it. thanks!